# Evil twins are not that evil: Qualitative insights into machine-generated prompts

### v1

@@ Acknowledgments
We thank Emmanuel Chemla, Mor Geva and audiences at SISA, CIMeC and at the online HiTZ seminar series for feedback. Our work was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101019291). This paper reflects the authors’ view only, and the funding agency is not responsible for any use that may be made of the information it contains.
Appendix A Token replacement examples
@@
We show randomly picked examples of single-token autoprompt replacements that do not affect the continuation, have a moderate effect on it or have a strong effect on it in tables 7, 8 and 9, respectively.
autoprompt: Processing<EOT> Launch/life},\$ Watson saw1949mL bigger wing
continuation: , a new engine, and a new propeller.
autoprompt: really dwarfs/blackados haben send extraordinarily overwhelmingly excessive\$]{} abundance
continuation: of heavy elements in their atmospheres.
autoprompt: approachò keep**_ mystery,. novel reportedly/council_** enjoys
continuation: keeping the reader in suspense.
autoprompt: impressive character<EOT> galactic Avengers drops<EOT><EOT>/further comics collected
continuation: in the Marvel Cinematic Universe.
autoprompt: champ241<EOT> GE 1870ista“ Japanese/rick dance art
continuation: form that was popular in the late 19th century.
@@
autoprompt: cancer<EOT> están<EOT> Card/Allenropical frig Jamaica describes humid
original continuation: tropical climate of the Caribbean.
modified continuation: tropical climate of the island of Jamaica.
modified continuation BLEU: 0.36
autoprompt: hired locals budget,** climbing destinations PullTown/LR oldest especially
original continuation: popular destination for climbers.
modified continuation: popular with climbers.
modified continuation BLEU: 0.23
autoprompt: schenken clergy?? KosovoABA<EOT> pledge regarding/loop your constant
original continuation: support of the Albanian Orthodox Church.
modified continuation: support for the Albanian Orthodox Church in Kosovo.
modified continuation BLEU: 0.43
autoprompt: <EOT>ITAL<EOT>/Angelesño<EOT> Denote 0̆415perorachusetts as
original continuation: the capital of the United States.
modified continuation: the capital of the United States of America.
modified continuation BLEU: 0.61
autoprompt: everyoneDaily tracking/idea>{{Self calendar??Its…exceedingly
original continuation: difficult to keep track of everything.
modified continuation: difficult to keep track of all the things that I want to do.
modified continuation BLEU: 0.28
@@
autoprompt: laccyt<EOT>ALTHICAN Brown jazz<EOT>/STan indispensable
original continuation: part of the American folk tradition.
modified continuation: to the development of the American style of jazz.
modified continuation BLEU: 0.11
autoprompt: Off"arn careers Birmingham lion2005 ballet/mediated Barry starred
original continuation: in the West End production of The Lion King.
modified continuation: in the film, which was released in the United States in 2005.
@@
autoprompt: Interview’————“ Heisenberg masterpiece/poverty Summer Fire books brand
original continuation: new introduction by the author.
modified continuation: new.
modified continuation BLEU: 0.02
autoprompt: tonnes Catholics/i Which Esc<EOT> have syn survived many factions
original continuation: including the Roman Catholic Church.
modified continuation: and are still in use.
modified continuation BLEU: 0.04
autoprompt: >] Publishingigenous Cemetery Once Anventh century/losing coffin had
original continuation: been buried in a pagan burial ground.
modified continuation: been found in the woods, the family decided to bury it in the family plot.
modified continuation BLEU: 0.01
@@
Appendix B Results with other models
B.1 Data-set statistics
Pythia-6.9B
As it is very time-consuming to extract autoprompts for this larger model, we limited the data-set to 208 entries. The average original prompt length is of 39.3 tokens (s.d. 13.4); that of the continuations is of 8.4 tokens (s.d. 2.4).
OLMo-1B
The data-set contains 500 entries. The average original prompt length is of 38.4 tokens (s.d. 11.2); that of the continuations is of 8.5 tokens.
B.2 Pruning autoprompts
  • Proportion of prunable autoprompts and average (s.d.) tokens pruned:
    Pythia-6.9B
    73.2% of autoprompts are pruned, with 2.6 (s.d. 1.6) tokens removed on average.
    OLMo-1B
    60.0% of autoprompts are pruned, with 1.9 (s.d. 1.1) tokens removed on average.
  • Token pruning distribution by position is shown in Fig. 6 (left: Pythia-6.9B; right: OLMo-1B).
Figure 6: Counts of pruned (dark orange) and kept (yellow) tokens in the autoprompts by position, in Pythia-6.9B (left) and OLMo-1B (right).
B.3 Replacing autoprompt tokens
For OLMo, we estimate the top 10k most frequent tokens to be used in the replacement experiments using a sample of approximately 10 billion tokens from the Dolma corpus, which was used to train this model. (Soldaini et al., 2024).
Proportions of replacement effect type by position are reported in Figure 7 (left: Pythia-6.9B; right: OLMo-1B).
Figure 7: Average proportions of replacement effect types in the autoprompts by position, aligned from right for Pythia-6.9B (left) and OLMo-1B (right) (whiskers show standard deviations). Null-effect replacements leave the continuation unchanged. Moderate replacements have BLEU of at least 0.2. Strong replacements have BLEU below 0.2.
Figure 8: Counts of pruned (dark orange) and kept (yellow) tokens in the original prompts, by position, in Pythia-6.9B (left) and OLMo-1B (right).
B.4 Shuffling autoprompt tokens
Average BLEU (s.d.) when shuffling all tokens vs. keeping last token fixed:
Pythia-6.9B
shuffling all tokens: 0.03 (s.d. 0.04); keeping last fixed: 0.06 (s.d. 0.11); paired t-test significant at $p<0.001$ (also if last-fixed is compared to random-non-last-fixed).
OLMo-1B
shuffling all tokens: 0.02 (s.d. 0.01); keeping last fixed: 0.04 (s.d. 0.04); t-test significant at $p<0.001$ (also if last-fixed is compared to random-non-last-fixed).
B.5 Making prompts more autoprompt-like
Pruning
  • Proportion of prunable prompts and average (s.d.) tokens pruned:
    Pythia-6.9B
    99.5% of the original prompts are pruned, and the average number of pruned tokens is 23.8 (s.d. 13.2); 95.7% of the pruned prompts have at least one pruned token among the last 10.
    OLMo-1B
    100% of the original prompts are pruned, and the average number of pruned tokens is 23.4 (s.d. 12.3); 97% of the pruned prompts have at least one pruned token among the last 10.
  • Token pruning by position is reported in Figure 8 (left: Pythia-6.9B; right: OLMo-1B).
Replacement
Proportions of replacement effect type by position are reported in Figure 9 (left: Pythia-6.9B; right: OLMo-1B).
Figure 9: Average proportions of replacement effect types in the original prompts by position, aligned from right (whiskers show standard deviations) for Pythia-6.9B (left) and OLMo-1B (right). Null-effect replacements leave the continuation unchanged. Moderate replacements have BLEU of at least 0.2. Strong replacements have BLEU below 0.2.
Shuffling
Average BLEU (s.d.) when shuffling all tokens vs. keeping last token fixed:
Pythia-6.9B
shuffling all tokens: 0.02 (s.d. 0.02); keeping last fixed: 0.03 (s.d. 0.05); paired t-test significant at $p<0.001$ (also if last-fixed is compared to random-non-last-fixed).
OLMo-1B
shuffling all tokens: 0.02 (s.d. 0.03); keeping last fixed: 0.03 (s.d. 0.05); paired t-test significant at $p<0.001$ (also if last-fixed is compared to random-non-last-fixed).
Appendix C Computing resources
@@
All experiments were run on a cluster composed of 11 nodes with 5 NVIDIA A30 GPUs each. The autoprompt search for Pythia-1.4B took approximately 600 GPU hours. Pruning, replacement and shuffling experiments for Pythia-1.4B took 1500 GPU hours overall. Compute demand for the other models was comparable.
Appendix D Assets
@@
AI use disclosure: we used Copilot and ChatGPT for assistance in code writing and in manuscript typesetting.

### v3

@@ Acknowledgments
We thank Emmanuel Chemla, Mor Geva and audiences at SISA, CIMeC and at the online HiTZ seminar series for feedback. Our work was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101019291). We also received funding from the Catalan government (AGAUR grant SGR 2021 00470). This paper reflects the authors’ view only, and the funding agencies are not responsible for any use that may be made of the information it contains.
Appendix A Results with other models
Data-set statistics
Following the procedure described in Section 3, we build a data-set for each of the following models: Pythia-1.4B, OLMo-1B, OLMo-7B, OLMo-7B-Instruct. Data-set statistics are presented in Table 7.
Model Data-set size Original prompt length Continuation length
Pythia-1.4B 2473 38.6 (11.7) 9.4 (2.7)
Pythia-12B 129 37.5 (11.0) 7.9 (1.7)
OLMo-1B 500 38.4 (11.2) 8.5 (2.0)
OLMo-7B 115 38.9 (10.9) 8.3 (1.9)
OLMo-7B-Instruct 104 39.8(12.7) 8.4(2.7)
Table 7: The number of entries, average original prompt length (s.d.), and average continuation length (s.d.) of the additional data-sets.
Pruning and shuffling
The results of the autoprompt pruning and shuffling experiments are presented in Table 8. For all models, there is a difference in BLEU when shuffling all tokens vs. keeping last token fixed (paired t-test significant at p < 0.01). Token pruning distribution by position is shown in Figures 17 18 19 20 21 (a).
Model Pruning rate
Tokens
pruned
BLEU
shuffle all
BLEU
shuffle except last
Pythia-1.4B 60% 1.2 (1.3) 0.03 (0.03) 0.05 (0.07)
Pythia-12B 86% 2.9 (1.9) 0.04 (0.05) 0.09 (0.12)
OLMo-1B 60% 1.2 (1.3) 0.02 (0.01) 0.04 (0.04)
OLMo-7B 60% 1.2 (1.4) 0.02 (0.02) 0.04 (0.05)
OLMo-7B-Instruct 27% 0.4(0.7) 0.01 (0.01) 0.02 (0.03)
Table 8: Results of the pruning and shuffling experiments of autoprompts. Pruning rate is the proportion of prompts in which at least one token could be pruned. There is a difference between BLEU scores when shuffling all tokens vs. keeping last token fixed for all models (paired t-test significant at p < 0.01).
Replacement
For OLMo models, we estimate the top 10k most frequent tokens to be used in the replacement experiments using a sample of approximately 10 billion tokens from the Dolma corpus, which was used to train this model. (Soldaini et al., 2024). Proportions of replacement effect type by position are reported in Figures 17 18 19 20 21 (b).
Semantic consistency
Similarly, we measure the semantic consistency of the equivalent sets in Figures 17 18 19 20 21 (c).
A.1 Making prompts more autoprompt-like
Model Pruning rate
Pruning rate
last 10 tokens
Tokens
pruned
BLEU
shuffle all
BLEU
shuffle except last
Pythia-1.4B 99% 95% 21.9(12.3) 0.02 (0.03) 0.03 (0.05)
Pythia-12B 100% 99% 23.8(11.7) 0.03 (0.04) 0.05 (0.12)
OLMo-1B 100% 97% 23.4(12.3) 0.02 (0.03) 0.03 (0.05)
OLMo-7B 100% 98% 24.0(12.6) 0.02 (0.02) 0.04 (0.06)
OLMo-7B-Instruct 92% 90% 21.4(14.6) 0.01 (0.01) 0.02 (0.03)
Table 9: Results of the pruning and shuffling experiments of original prompts. There is a difference between BLEU scores when shuffling all tokens vs. keeping last token fixed for all models except OLMo-7B-Instruct (paired t-test significant at p < 0.01).
Pruning and shuffling
The results of the pruning and shuffling experiments of natural prompts are presented in Table 9. For all models except OLMo-7B-Instruct, there is a difference in BLEU when shuffling all tokens vs. keeping last token fixed (paired t-test significant at p < 0.01). Token pruning by position is reported in Figures 17 18 19 20 21 (a).
Replacement
Proportions of replacement effect type by position are reported in Figures 17 18 19 20 21 (b).
Semantic consistency
Semantic consistency of the equivalent sets are reported in Figures 17 18 19 20 21 (c).
Appendix B Token replacement examples
@@
We show randomly picked examples of single-token autoprompt replacements that do not affect the continuation, have a moderate effect on it or have a strong effect on it in tables 10, 11 and 12, respectively. In addition, Table B provides additional examples of replacement with null-effect, where tokens are colored based on the number of equivalent substitutes.
autoprompt: crossesFootnote Several/Accordingly panels accidentally have feather 517 chant collectively continuation: , as if they were a single bird. autoprompt: <EOT><EOT> unbelievable/lands deep intuitive continuation: understanding of the human condition. autoprompt: throne wereivered/printceryassociated pursuing Somerset where continuation: he was arrested and imprisoned in the Tower of London. autoprompt: tables/app Its 590 chapel wide Rosadian continuation: marble floor is the largest in the world. autoprompt: FLOAT depicts 1933}}?significant professional MMA classic/adapt pickup after continuation: a long absence from the sport.
@@
autoprompt: TLR/ract<EOT> pledmanifest behalf breached adiabarfollow
original continuation: the terms of the settlement agreement.
modified continuation: ed the terms of the agreement.
modified continuation BLEU: 0.49
autoprompt: stylish 1876 painting escaped creditors Anne/Im decides
original continuation: to sell it to pay off her debts.
modified continuation: to sell it to a wealthy collector.
modified continuation BLEU: 0.32
autoprompt: Round/22 transferring’\$, when splash energy
original continuation: is transferred to the surrounding water.
modified continuation: is transferred to the surface of the liquid.
modified continuation BLEU: 0.32
autoprompt: <EOT><EOT>/Two<EOT> Return onto Broadway where Gardens Airport coincide
original continuation: with the Garden State Parkway.
modified continuation: with the intersection of Broadway and the Garden State Parkway.
modified continuation BLEU: 0.45
autoprompt: <EOT>/Where Treaty<EOT>Unionduring
original continuation: the American Revolutionary War.
modified continuation: the Revolutionary War.
modified continuation BLEU: 0.33
@@
autoprompt: FIRST<EOT>protective talentsSarah/bias<EOT> uses learning technical
original continuation: skills to protect herself and others.
modified continuation: skills to protect the user from harm.
modified continuation BLEU: 0.15
autoprompt: fiveZero Parl/insterson homicide By 1867 Provincialswick warranted
original continuation: the establishment of a police force.
modified continuation: the appointment of a police magistrate.
modified continuation BLEU: 0.19
autoprompt: stylish 1876 painting escaped creditors/() Anne decides
original continuation: to sell it to pay off her debts.
modified continuation: to take a break from her work and go for a walk.
modified continuation BLEU: 0.02
autoprompt: Pretty.“byterBlood "A realistic work/convicted compatible
original continuation: with the other works of the author
modified continuation: with the "Atheist" and "Atheist" "Atheist" an[…]
modified continuation BLEU: 0.02
autoprompt: ——————————————————terrorismworker is/exit killed
original continuation: in a suicide bombing in Iraq.
modified continuation: in Iraq
@@
Autoprompts Generated Continuation
ipzig suspended Emma Leigh ’s 956 time commitments and as …a result, she was able to devote herself to her writing.
[…]
Emma: ’ women’, ’ woman’, ’ tau’, ’ wife’, ’ sex’, ’ da’, ’ ura’, ’ snow’, …
commitments: ’ attention’, ’ traditional’, ’ burden’, ’ politics’, ’ busy’, …
and: ’ how’, ’ and’, ’ et’, ’ that’, ’ ad’, ’ und’, …
as: ’ as’
suspended: 0 substitutes
COMMAND uminous photometry has ? continually undermined …the idea that the Sun is a star.
[…]
undermined: ’ rejected’
photometry: 0 substitutes.
Original Prompt Generated Continuation
South Australian with Wales requested that Major General William ois be appointed to command the expedition.
[…]
Major: ’ sub’,’ major’,’ officer’,’ Sub’
General: ’ general’,’ officer’,’ Gen’,’ Captain’
that: 0 substitutes.
erected in 1896 listed of the sand stone market building , which was built in the late 19th century.
the: ’ an’,’ its’,’ some’,’ great’,’ old’,’ further’,’ additional’,’ especially’, …
stone: ’aligned’,’ized’,’á’,’ette’,’ished’,’ stone’,’split’,’white’,’ulate’, …
building: ’ court’,’ structure’,’ complex’,’ House’,’ unit’,’ track’,’ society’,’ mechanism’, …
Table 13: Additional Examples of replacement for autoprompts and original prompts (10 last tokens). Color represents the number of substitutes: $>50$ substitutes and $<50$ substitutes. When there are less than 50 substitutes, a random subset is displayed. Difficult-to-render characters are replaced by ’?’.
Large substitution sets
Table 14 presents a larger list of tokens that are the most frequently found in large equivalent substitution sets. It mostly consists of subwords (sometimes corresponding to interpretable morphemes), digits and (fragments of) named entities, that do not carry strong semantic information. We conjecture that digits appear in cases where the exact number is not crucial for generating the continuation, but the notion of quantity or date is important (e.g., in that case $5$, $10$ or $18$ are equivalents). A similar interpretation holds for named entities. For instance, ‘Moore’, ‘Brown’ or ‘Smith’ are common proper nouns that could refer to a typical North-American entity in an exchangeable way.
105 <EOT> () 113 ini Mar ran 25 be Green 101 56 ¸ mat ato ator Rock 35 115 Moore 114 SS lim eter ester abe uns Head Char ? ? ? ac aa aur ? ong ost den FF 54 a ab ud sl rim st ? 022 1 37 91 INT op ore ef rel cha ast ives 2000 ash arch ath eps ced Har fall met ales 140 Ros 5 48 21 33 101 38 75 600 ? ** oth aff ay br LE ? anch der cy ris fer inner ots arn Sil Red ros , cent Brown ister dec 50 44 40 28 500 ? 41 46 85 112 93 45 57 n y em ard na eng ater ree EE urg ? ? ord oto 1 - ign sen ans ach dis ush Br Cam pol one ? ner 18 42 48 250 96 Bell 94 117 19 ial pp ba star ? mar har hor Tor lan vi Sl 1986 arg ender Sand First Po Che Jones 100 per ative ways hal ben bon Ross 6 ? 30 46 31 65 400 1000 LE 33 am Bar ace amb Ab ull Reg ena orm stra s co ards win iver Ha ? pin init dist pres Ray 47 Smith 120 128 number án Kar 4 70 26 27 98 200 80 200 2000 45 500 49 110 63 13 el ale za ute
Table 14: List of the $240$ tokens most frequently found in large equivalent sets (observed in about $10\%$ of the sets). Difficult-to-render characters are replaced by ’?’. Tokens are marked as follows: digit, artifact and punctuation, named entity, subword and morphological unit.
Appendix C Human annotation of token importance
The four authors of this paper manually annotated the autoprompt dataset (for the Pythia 6.9B model), assigning a binary label to each token to indicate whether it is intuitively important for generating the continuation. The annotation process was conducted a posteriori, with annotators having access to the full autoprompt and its corresponding generated continuation. To minimize order effects, tokens were randomized before annotation. As an illustration, Table 15 provides some annotated examples.
In total, $1971$ tokens have been labeled by at least one annotator. We assess inter-annotator agreement on a subset of $50$ tokens: the four annotators agreed on $63.3\%$ of the examples, and at least 3 annotators agreed on $89.8\%$ of the examples.
Table 16 provides the full comparison between importance annotation and equivalent set size.
Autoprompt Continuation Token to annotate Annotation <EOT><EOT><EOT><EOT> _____<EOT> European devastated 1941 prompted …the United States to enter the war. <EOT> 0 clut British’); from 40 UTC anyone emitted one speaker …in the UK would be heard in the UK. speaker 1 Croat trainersers?? Greece Catholics Deltaclaim independently …from the Greek Orthodox Church. Greece 1 Shan Marxist Augustine Fran??ois State TERis Railroad serves ? …the city of Terre Haute, Indiana. Railroad 1 defenseLittle<EOT>Content<EOT> forests constitute reliable leaving habitats …for many species of birds and mammals. leaving 1 earliest 2??–¢ flares when Jacksonville dominancethe metropolitan …area was still in its infancy. dominance 0 ‡?Å Berry paternalEffects cephal???. Johann Born evenublished …a book on the subject in 1775. Johann 1 1953 that Dix Teachers 1989 Archbishopreceived a Hamilton Legacy …Award for his contributions to education. that 0
Table 15: We provide 8 examples of human annotations of token’s importance. Using the autoprompt and the continuation, the annotator had to label the “token to annotate” as important (labeled with 1) or not important (labeled with 0). Difficult-to-render characters are replaced by ’?’
All Lang-like Non-Linguistic
Imp. Not-Imp. Imp. Not-Imp. Imp. Not-Imp.
Average $163.4$ $372.2$ $150.8$ $298.7$ $268.6$ $538.4$
Median $2.0$ $35.0$ $2.0$ $23.0$ $13.5$ $83.0$
Table 16: Comparison of the equivalent set size (average and median) between tokens classified as Important (Imp.) or Not-Important (Not-Imp.) by human experts, across different token categories.
Appendix D Computing resources
@@
All experiments were run on a cluster composed of 11 nodes with 5 NVIDIA A30 GPUs each. The autoprompt search for Pythia-1.4B took approximately 600 GPU hours. Pruning, replacement and shuffling experiments for Pythia-1.4B took 1500 GPU hours overall. Compute demand for the other models was comparable, although we had to settle for smaller autoprompt sets.
Appendix E Assets
@@
(a) Count of pruned (orange) and kept (yellow) tokens (left: autoprompt; right: original prompt).
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original).
(c) Semantic similarity between substitutes (left: autoprompt; right: original prompt).
Table 17: Pythia-1.3B: reproducing pruning and replacement experiments.
(a) Count of pruned (orange) and kept (yellow) tokens (left: autoprompt; right: original prompt).
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original).
(c) Semantic similarity between substitutes (left: autoprompt; right: original prompt).
Table 18: OLMo-1B: reproducing pruning and replacement experiments.
(a) Count of pruned (orange) and kept (yellow) tokens (left: autoprompt; right: original prompt).
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original).
(c) Semantic similarity between substitutes (left: autoprompt; right: original prompt).
Table 19: OLMo-7B: reproducing pruning and replacement experiments.
(a) Count of pruned (orange) and kept (yellow) tokens (left: autoprompt; right: original prompt).
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original).
(c) Semantic similarity between substitutes (left: autoprompt; right: original prompt).
Table 20: Pythia-12B: reproducing pruning and replacement experiments.
(a) Count of pruned (orange) and kept (yellow) tokens (left: autoprompt; right: original prompt).
(b) Average proportions of replacement effect types by position (left: autoprompt; right: original).
(c) Semantic similarity between substitutes (left: autoprompt; right: original prompt).
Table 21: OLMo-7B-Instruct: reproducing pruning and replacement experiments.